# Site-selective renormalization and competing magnetic instabilities in paramagnet ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$

### 18 Mar 2026 v1

@@ Abstract
Quantum spin liquids (QSLs) are exotic phases of matter characterized by long-range entanglement and the absence of magnetic order even at zero temperature. Here, we present a comprehensive theoretical study of the frustrated magnet ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ to elucidate its electronic and magnetic properties. We uncover two inequivalent Cu sites with fundamentally distinct oxygen coordination—octahedral CuO6 and axially compressed CuO8—that give rise to completely opposite crystal-field splittings. This inversion places the unpaired hole in the $d_{z^{2}}$ orbital at the Cu-2 site, while Cu-1 maintains conventional $d_{x^{2}-y^{2}}/d_{xy}$ character, producing narrow bands at commensurate filling. Dynamical mean-field theory (DMFT) calculation reveals a selective band-renormalization of orbitals from the two Cu ions with Cu-1 undergoing a Mott transition while Cu-2 remaining metallic under electronic correlations. Crucially, FLEX calculations demonstrate that multiple magnetic instabilities compete with nearly equal strength: the spin susceptibility lacks dominant peaks, and the leading eigenvalues approach unity simultaneously across all wavevectors with increasing interaction. This competitive interplay, originating from the distinct local environments and geometric frustration on the triangular lattice, agrees well with the absence of long-range magnetic order in experiment. Our results suggest ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ as a promising QSL candidate where the unique combination of disparate crystal-field environments, strong correlations, and competing exchange interactions conspire to stabilize an exotic quantum ground state.
@@ I Introduction
The pursuit of the quantum spin liquid (QSL) state—a highly entangled phase where magnetic moments remain dynamic at absolute zero—is a cornerstone of modern condensed matter physics [1]. Theoretically, the antiferromagnetic triangular lattice is the canonical platform for such a phase, where geometrical frustration precludes simple Néel order. While the pure $J_{1}$ isotropic Heisenberg model on a triangular lattice eventually settles into a 120 ordered state [2], the introduction of competing next-nearest-neighbor interactions ($J_{2}/J_{1}\sim(0.07,0.15))$ is predicted to stabilize exotic gapped or gapless QSL phases [3].
Despite decades of experimental searching, a “firmly accepted” triangular QSL remains elusive due to persistent material-specific challenges. First, structural disorder, such as the Mg/Ga site-mixing in the widely studied YbMgGaO4, can mimic QSL signatures by inducing a random singlet or spin-glass state rather than a true QSL phase [4, 5]. Second, many 2D candidates suffer from subtle interlayer couplings or exchange anisotropies that drive the system toward long-range order at sub-Kelvin temperatures, as seen in various delafossite-related structures [6].
In the last decade, the experimental focus has shifted from disorder-prone materials toward “cleaner” triangular systems. Significant progress has been made with the alkali-based delafossites $AYbCh_{2}$ (AA = Na, K; $Ch$ = O, S, Se). Most notably, NaYbSe2 has emerged as a premier QSL candidate due to its lack of long-range magnetic ordering and the presence of coherent, fractionalized spinon excitations. Early studies proposed a gapless spinon Fermi surface ground state [7, 8, 9] from ultralow temperature susceptibility, specific heat, and thermal conductivity measurements. Similarly, TlYbSe2 has emerged as a rare example of a field-tunable QSL where gauge field excitations may play a critical role [10]. Furthermore, the discovery of PrMgAl11O19 has introduced the possibility of an Ising-type QSL, expanding the search beyond isotropic Heisenberg symmetries [11, 12]. The “Ising-type” designation is firmly established across all studies. Magnetic susceptibility measurements consistently show that the c-axis is the easy axis, with in-plane (a-b plane) susceptibility being approximately two orders of magnitude smaller [13, 14, 15, 16]. This extreme anisotropy, combined with the perfect triangular lattice formed by Pr3+ ions, makes PrMgAl11O19 a rare realization of the triangular-lattice Ising antiferromagnet (TLIAF) model. While all studies of this system agree on the Ising anisotropy and the absence of magnetic order, there is currently an active debate regarding whether the QSL ground state is gapless or gapped [15, 17, 16]. Among the triangular systems, there is also a class of materials named cluster Mott insulators, where molecular units rather than individual atoms carry the effective spin-1/2 moments. The ${\mathrm{Mo}}_{3}{\mathrm{O}}_{8}$-based kagome systems (${\mathrm{Li}}{\mathrm{Zn}}_{2}{\mathrm{Mo}}_{3}{\mathrm{O}}_{8}$ family) [18, 19, 20, 21, 22, 23, 24, 25], and Nb3X8 family (X = Cl, Br, I) [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] exemplify this class, where trimerization leads to effective triangular lattices of spin clusters. The $1T\text{-}{\mathrm{TaS}}_{2}$ system similarly features star-of-David clusters forming a triangular arrangement  [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. These molecular QSL candidates bridge the gap between atomic-scale frustrated magnets and emergent lattice geometries.
@@
In this work, we present a theoretical investigation of the spin-$1/2$ three-dimensional frustrated magnet ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$. By combining density-functional theory (DFT) calculations, perturbative many-body methods, and dynamical mean-field theory (DMFT), we analyze the electronic and magnetic responses of this material, offering insight into its unusual low-energy excitations. The non-magnetic electronic structure is governed by two inequivalent Cu octahedra that exhibit opposite trigonal distortion modes, leading to fundamentally different crystal-field splittings. The low-energy degrees of freedom are described by narrow bands at commensurate electron filling, which undergo a Mott transition under the influence of strong electronic correlations. Calculations of the spin susceptibility and LSDA+$U$ total energies for various magnetic configurations reveal strong competition among multiple magnetic wave vectors, suggesting that the ground state may realize a quantum spin liquid.
The remainder of this paper is organized as follows. In Section II, we first discuss the electronic structure obtained from DFT and DFT+DMFT, demonstrating the unique crystal-field of the two Cu ions and electronic band renormalization. We then examine the magnetic couplings using two complementary approaches. First, we analyze the paramagnetic instability of the spin susceptibility within the fluctuation-exchange approximation (FLEX). Second, we compare the total energies of various magnetic configurations, confirming their competitive nature and the absence of a dominant magnetic instability. Finally, in Section III, we summarize our conclusions and discuss their implications for experimental identification of a quantum spin liquid state in ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$.
@@ II Results
${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ crystallizes in the $R\bar{3}m$ structure [51], illustrated in Fig. 1(a). The three-dimensional framework features two inequivalent Cu2+ ions, each forming edge-shared triangular lattices. This geometry gives rise to strong magnetic frustration and deviates from conventional ordered magnetic arrangements. The structure comprises interconnected polyhedra of copper and antimony coordinated by oxygen, with yttrium cations occupying interstitial sites to maintain charge balance. The presence of two distinct copper environments introduces competing magnetic interactions that inhibit long-range magnetic order and may stabilize exotic quantum ground states. This structural complexity establishes ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ as a promising platform for investigating quantum spin liquid behavior and related emergent phenomena in low-dimensional magnets. In the following we will illuminate the consequence of the two different Cu local environment, which uniquely characterizes ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$ from other candidate materials.
Figure 1: Crystal and electronic structure of Y3Cu2Sb3O14. (a) The crystal structure of ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$, in which the two inequivalent copper sites (yellow atoms), Cu-1 and Cu-2, individually form triangular planes. (b) shows the two different local environment of Cu-1 and Cu-2. In Cu-2, an additional shorter Cu-O bond along z axis inverts the crystal field created by the regular trigonal distortion in Cu-1. (c) The DFT electronic band structure of the primitive cell is shown along high-symmetry paths in the first Brillouin zone. The orange line around the Fermi level denotes the three bands with their Wannier functions shown in (d). Additionally including the purple bands leads to a ten bands effective model. Both the three-band and the ten-band models show excellent agreement with the DFT results. (d) The projected Wannier functions of the three yellow bands indicate that, among the three orbitals, two orbitals are from the Cu-1 ion and one from Cu-2 ion.
@@ II.1 Inverted crystal-field splittings
The two inequivalent Cu sites experience markedly different oxygen coordination environments: $\mathrm{CuO}_{6}$ and $\mathrm{CuO}_{8}$. Each Cu sublattice forms an in-plane triangular lattice with a layered geometry. For clarity, we denote the Cu atom at the corner of the unit cell $(0,0,0)$ as Cu-1 and the one at the center $(1/2,1/2,1/2)$ as Cu-2. At the Cu-1 site, each Cu atom is coordinated by six oxygen atoms with equal bond lengths of 1.95 Å, but the bond angles deviate significantly from $90^{\circ}$, resulting in a compressive trigonal distortion. This coordination departs from a regular octahedral environment and instead exhibits local symmetry close to $D_{3d}$. Under such a crystal field, the fivefold-degenerate $d$ orbitals split into three distinct manifolds: a nondegenerate $A_{1g}$ orbital (predominantly $d_{z^{2}}$) and two sets of doubly degenerate $E_{g}$ levels, with the typical energy ordering
@@
The absence of bond-length variations suggests that the Jahn-Teller distortion expected for a $d^{9}$ (Cu2+) ion is effectively suppressed, likely due to lattice constraints or covalent bonding effects.
@@ II.2 DFT electronic structure
Having established the local valence configuration and crystal field, we now examine the Bloch band structure of ${\mathrm{Y}}_{3}{\mathrm{Cu}}_{2}{\mathrm{Sb}}_{3}{\mathrm{O}}_{14}$. Figure 1(c) shows the electronic dispersion along a chosen high-symmetry path. DFT predicts a metallic ground state, with Cu-$d$ orbitals dominating the low-energy states near the Fermi level. The projected density of states (PDOS) in the right panel of Fig. 1(c) reveals distinct $d$-orbital contributions from Cu-1 and Cu-2. Consistent with the crystal-field analysis above, the two sites share the same local $D_{3d}$ point-group symmetry, but the shorter axial Cu–O bonds at Cu-2 induce a complete inversion of the $d$-orbital energy ordering relative to Cu-1. At Cu-1, the $d_{z^{2}}$ orbital lies at approximately 2-2 eV as the lowest among the five $d$ states, whereas at Cu-2 it becomes the highest. Consequently, Cu-1 contributes a twofold $E_{g}$ manifold, while Cu-2 provides a nondegenerate $A_{1g}$ state, suggesting that the minimal low-energy model must include at least three orbitals.
Using maximally localized Wannier functions as implemented in Wannier90 [53], we construct a three-orbital model (yellow solid line in Fig. 1(c)) that accurately reproduces the DFT bands. The corresponding Wannier functions, displayed in Fig. 1(d), exhibit two distinct symmetry types. Two of the Wannier functions reside on Cu-1 and reflect the twofold $E_{g}$ symmetry, while the third is centered on Cu-2 and displays $C_{3}$-symmetric $A_{1g}$ character, fully consistent with the two distinct $D_{3d}$ crystal-field splittings. Three $d$ valence electrons occupy the two $E_{g}$ Wannier functions at Cu-1 while the $A_{1g}$ Wannier function at Cu-2 is occupied by one electron. A more complete description includes all five $d$ orbitals from each Cu ion, resulting in a ten-orbital model (blue solid line in Fig. 1(c)). States derived from Y, Sb, and O lie at higher binding energies and do not contribute to the low-energy effective model.
The conclusion that the Cu-1 and Cu-2 $d$ orbitals are integer-filled holds only in the atomic limit. Upon forming Bloch wavefunctions—particularly after downfolding to effective low-energy models—the orbital filling often deviates substantially from integer values. Nonetheless, our DFT calculations reveal that the electron occupation of the twofold degenerate $E_{g}$ bands at the Cu-1 site remains exactly 3, whereas the $A_{1g}$ state at the Cu-2 site is occupied by approximately 1.15 electrons. This deviation likely arises from stronger Cu–O hybridization associated with an additional shorter Cu–O bond (see Fig. 1(b)).
@@ II.3 Correlated Electronic Structure
Given the distinct filling levels of the two Cu ions and the narrow bandwidth near the Fermi level, electronic correlations are expected to play a significant role in shaping the electronic states. In this section, we investigate the evolution of the electronic structure under the influence of Coulomb interactions. To this end, we employ the DFT+DMFT scheme implemented in our in-house “Package for Analyzing Correlated Systems” (PACS) [54, 55] and cross-validate our results using the TRIQS package [56, 57]. In both implementations, we restrict the correlated subspace to the three partially filled bands near the Fermi level (indicated by yellow solid lines in Fig. 1(d)).
@@
Since DMFT is a local approximation, careful consideration must be given to the construction of the impurity problem. The two Cu sites experience distinct local crystal fields. Among the three bands near the Fermi level, two originate from Cu-1 and one from Cu-2. Including all three Wannier functions in a single impurity problem would mix the two independent Cu ions via electronic correlations, potentially obscuring their distinct physical characters. A more appropriate approach is to define two separate impurity problems, one for each Cu site. In the following, we examine both schemes and compare their outcomes in Fig. 2.
@@
In the first scheme, we treat all three Wannier functions within a single impurity problem. In addition to the intra-site couplings between the two degenerate orbitals on Cu-1, the $A_{1g}$ state on Cu-2 is coupled to them via single-particle hybridizations and density–density Coulomb interactions, resulting in an effective three-orbital impurity model. We solved this model for two interaction strengths, $U=1.0$ and $2.0$ eV, with $\beta=30$ and $J=0.1$ eV. As shown in Fig. 2(a), the resulting correlated spectral function exhibits broad features spanning from 1.5-1.5 eV to $1.0$ eV. The imaginary parts of the self-energies for the two Cu sites show distinct behaviors. The two $E_{g}$ Wannier states on Cu-1 remain nearly degenerate (red line with circles) and exhibit a smaller self-energy than that of the $A_{1g}$ state on Cu-2 (blue line with squares), indicating weaker band renormalization for Cu-1. However, increasing $U$ to $2.0$ eV lifts the degeneracy of the $E_{g}$ states, resulting in two distinct self-energies for Cu-1. In this unified impurity treatment, electronic correlations enhance the single-particle hybridization among the three Wannier states, leading to an unphysical breaking of the $E_{g}$ symmetry.
@@
A more physically sound DMFT framework assigns the three Wannier states to separate impurity problems based on their spatial localization. In the second scheme, we therefore treat the two $E_{g}$ states on Cu-1 and the $A_{1g}$ state on Cu-2 in two independent impurity models, solving for their respective self-energies separately. Both self-energies are incorporated simultaneously into the DMFT self-consistency loop and the DFT+DMFT upfolding/downfolding procedure. The resulting spectral functions and self-energies are presented in Fig. 2(c) and (d). First, similar to the unified treatment, the system remains metallic across the parameter range studied. However, the self-energy now preserves the expected symmetry: the $E_{g}$ states remain nearly degenerate and distinct from the $A_{1g}$ state at both interaction strengths. Second, we observe that the integer-filled $E_{g}$ states are more sensitive to electronic correlations. As $U$ increases from 1 to 2 eV, the self-energy of the $E_{g}$ states evolves from metallic toward more insulating behavior, while that of the $A_{1g}$ state retains its metallic character. This stark contrast arises from the difference in orbital fillings: the integer-filled $E_{g}$ states on Cu-1 experience stronger band renormalization, whereas the partially filled $A_{1g}$ state remains relatively weakly correlated. Consequently, $\mathrm{Y}_{3}\mathrm{Cu}_{2}\mathrm{Sb}_{3}\mathrm{O}_{14}$ exhibits site-selective band renormalization, where bands derived from Cu-1 orbitals are more strongly correlated than those from Cu-2. The observation of site-selective band renormalization in $\mathrm{Y}_{3}\mathrm{Cu}_{2}\mathrm{Sb}_{3}\mathrm{O}_{14}$ offers a promising explanation to the two-step spin condensation observed in experiment, where the charge degree of Cu-2 is less correlated and only freeze at much lower temperature, inhibiting the condensation of spins around Cu-2.
@@
This site-selective behavior is reminiscent of the extensively studied rare-earth nickelates RRNiO3 (with $R=$ Sm, Eu, Y, Lu), in which Ni occupies two distinct sublattices characterized by different Ni–O bond lengths [58]. In those materials, Ni ions in long-bond octahedra adopt a nominal $d^{8}$ configuration, whereas those in short-bond octahedra exhibit a $d^{8}\underline{L}^{2}$ configuration, where two ligand holes compensate two electrons in the Ni $e_{g}$ orbitals. This leads to inequivalent Ni sites and differential band renormalization under electronic correlations [59, 60, 61, 62]. In $\mathrm{Y}_{3}\mathrm{Cu}_{2}\mathrm{Sb}_{3}\mathrm{O}_{14}$, the situation is even more striking: the two Cu2+ ions experience completely inverted crystal fields, giving rise to distinct correlation responses from the two Cu sites.

### 19 Mar 2026 v2

# Site-selective renormalization and competing magnetic instabilities in paramagnet Y3Cu2Sb3O14

@@ Abstract
Quantum spin liquids (QSLs) are exotic phases of matter characterized by long-range entanglement and the absence of magnetic order even at zero temperature. Here, we present a comprehensive theoretical study of the frustrated magnet Y3Cu2Sb3O14 to elucidate its electronic and magnetic properties. We uncover completely opposite crystal-field splittings of the two inequivalent Cu sites owing to their fundamentally distinct oxygen coordination—trigonal distorted octahedral CuO6 and axially compressed CuO8. This inversion places the unpaired hole in the $d_{z^{2}}$ orbital at the Cu-2 site, while Cu-1 maintains conventional $d_{x^{2}-y^{2}}/d_{xy}$ character, which results in a selective band-renormalization of orbitals from the two Cu ions. We further find multiple magnetic instabilities competing with nearly equal strength in this system: the spin susceptibility lacks dominant peaks, and the leading eigenvalues approach unity simultaneously across all wavevectors with increasing interactions. This competitive interplay, originating from the distinct local environments and geometric frustration on the triangular lattice, agrees well with the absence of long-range magnetic order in experiment. Our results support Y3Cu2Sb3O14 as a promising QSL candidate where the unique combination of disparate crystal-field environments, strong correlations, and competing exchange interactions conspire to stabilize an exotic quantum ground state.
@@ I Introduction
The pursuit of the quantum spin liquid (QSL) state—a highly entangled phase where magnetic moments remain dynamic at absolute zero—is a cornerstone of modern condensed matter physics [1]. Theoretically, the antiferromagnetic triangular lattice is the canonical platform for such a phase, where geometrical frustration precludes simple Néel order. While the pure $J_{1}$ isotropic Heisenberg model on a triangular lattice eventually settles into a 120 ordered state [2], the introduction of competing next-nearest-neighbor interactions ($J_{2}/J_{1}\sim(0.06,0.17))$ is predicted to stabilize exotic gapped or gapless QSL phases [3, 4, 5].
@@
[…]
In this work, we present a theoretical investigation of the spin-$1/2$ three-dimensional frustrated magnet Y3Cu2Sb3O14. By combining density-functional theory (DFT) calculations, perturbative many-body methods, and dynamical mean-field theory (DMFT), we analyze the electronic and magnetic responses of this material, offering insight into its unusual low-energy excitations. The non-magnetic electronic structure is governed by two inequivalent Cu octahedra that exhibit opposite trigonal distortion modes, leading to fundamentally different crystal-field splittings. The low-energy degrees of freedom are described by narrow bands, which are renormalized by electronic correlations to different extent featuring a unique site-selective character. Calculations of the spin susceptibility reveal a strong competition among multiple magnetic wave vectors, suggesting that the ground state may realize a quantum spin liquid.
The remainder of this paper is organized as follows. In Section II, we first discuss the electronic structure obtained from DFT and DFT+DMFT, demonstrating the unique crystal-field of the two Cu ions and electronic band renormalization. We then examine the spin susceptibility within the fluctuation-exchange approximation (FLEX), confirming its competitive nature and the absence of a dominant magnetic instability. Finally, in Section III, we summarize our conclusions and discuss their implications for experimental identification of a quantum spin liquid state in Y3Cu2Sb3O14.
@@ II.1 Inverted crystal-field splittings
Figure 1: Crystal and electronic structure of Y3Cu2Sb3O14. (a) The crystal structure of Y3Cu2Sb3O14, in which the two inequivalent copper sites (yellow atoms), Cu-1 and Cu-2, individually form triangular planes. (b) shows the two different local environment of Cu-1 and Cu-2. In Cu-2, an additional shorter Cu-O bond along z axis inverts the crystal field created by the regular trigonal distortion in Cu-1. (c) The DFT electronic band structure of the primitive cell is shown along a high-symmetry path in the first BZ. The orange line around the Fermi level denotes the three bands with their Wannier functions shown in (d). Additionally including the purple bands leads to a ten bands effective model. Both the three-band and the ten-band models show excellent agreement with the DFT results. (d) The projected Wannier functions of the three yellow bands indicate that, among the three orbitals, two orbitals are from the Cu-1 ion and one from Cu-2 ion.
Y3Cu2Sb3O14 crystallizes in the $R\bar{3}m$ structure [56], illustrated in Fig. 1(a). The structure comprises interconnected polyhedra of copper and antimony coordinated by oxygen, with yttrium cations occupying interstitial sites to maintain charge balance. The three-dimensional framework features two inequivalent Cu2+ ions, each forming edge-shared triangular lattices. This geometry gives rise to strong magnetic frustration and deviates from conventional ordered magnetic arrangements.
The two inequivalent Cu sites experience markedly different oxygen coordination environments: CuO6 and CuO8. Each Cu sublattice forms an in-plane triangular lattice with a layered geometry. For clarity, we denote the Cu atom at the corner of the unit cell $(0,0,0)$ as Cu-1 and the one at the center $(1/2,1/2,1/2)$ as Cu-2. At the Cu-1 site, each Cu atom is coordinated by six oxygen atoms with equal bond lengths of 1.95 Å, but the bond angles deviate significantly from $90^{\circ}$, resulting in a compressive trigonal distortion. This coordination departs from a regular octahedral environment and instead exhibits local symmetry close to $D_{3d}$. Under such a crystal field, the fivefold-degenerate $d$ orbitals split into three distinct manifolds: a nondegenerate $A_{1g}$ orbital (predominantly $d_{z^{2}}$) and two sets of doubly degenerate $E_{g}$ levels, with the typical energy ordering
@@ II.2 DFT electronic structure
Having established the local valence configuration and crystal field, we now examine the Bloch band structure of Y3Cu2Sb3O14. Figure 1(c) shows the electronic dispersion along a chosen high-symmetry path in the Brillouin zone (BZ). DFT predicts a metallic ground state, with Cu-$d$ orbitals dominating the low-energy states near the Fermi level. The projected density of states (PDOS) in the right panel of Fig. 1(c) reveals distinct $d$-orbital contributions from Cu-1 and Cu-2. Consistent with the crystal-field analysis above, the two sites share the same local $D_{3d}$ point-group symmetry, but the shorter axial Cu–O bonds at Cu-2 induce a complete inversion of the $d$-orbital energy ordering relative to Cu-1. At Cu-1, the $d_{z^{2}}$ orbital lies at approximately 2-2 eV as the lowest among the five $d$ states, whereas at Cu-2 it becomes the highest. Consequently, Cu-1 contributes a twofold $E_{g}$ manifold, while Cu-2 provides a nondegenerate $A_{1g}$ state to the Fermi surface, suggesting that the minimal low-energy model must include at least three orbitals.
Using maximally localized Wannier functions as implemented in Wannier90 [58], we construct a three-orbital model (yellow solid line in Fig. 1(c)) that reasonably reproduces the DFT bands. The corresponding Wannier functions, displayed in Fig. 1(d), exhibit two distinct symmetry types. Two of the Wannier functions reside on Cu-1 and reflect the twofold $E_{g}$ symmetry, while the third is centered on Cu-2 and displays $C_{3}$-symmetric $A_{1g}$ character, fully consistent with the two distinct $D_{3d}$ crystal-field splittings. Three $d$ valence electrons occupy the two $E_{g}$ Wannier functions at Cu-1 while the $A_{1g}$ Wannier function at Cu-2 is occupied by one electron. A more complete description includes all five $d$ orbitals from each Cu ion, resulting in a ten-orbital model (blue solid line in Fig. 1(c)). States derived from Y, Sb, and O lie at higher binding energies and do not significantly contribute to the low-energy effective model.
The conclusion that the Cu-1 and Cu-2 $d$ orbitals are integer-filled holds only in the atomic limit. Upon forming Bloch wavefunctions—particularly after downfolding to effective low-energy models—the orbital filling often deviates substantially from integer values. Nonetheless, our DFT calculations reveal that the electron occupation of the twofold degenerate $E_{g}$ bands at the Cu-1 site remains 3, whereas the $A_{1g}$ state at the Cu-2 site is occupied by approximately 1.15 electrons. This deviation likely arises from stronger Cu–O hybridization associated with an additional shorter Cu–O bond (see Fig. 1(b)).
@@ II.3 Correlated Electronic Structure
Given the distinct filling levels of the two Cu ions and the narrow bandwidth near the Fermi level, electronic correlations are expected to behave differently in shaping the states of the electrons around the two Cu ions. In this section, we investigate the evolution of the electronic structure under the influence of Coulomb interactions. To this end, we employ the DFT+DMFT scheme implemented in our in-house “Package for Analyzing Correlated Systems” (PACS) [59, 60] and cross-validate our results using the TRIQS package [61, 62]. In both implementations, we restrict the correlated subspace to the three partially filled bands near the Fermi level (indicated by yellow solid lines in Fig. 1(d)).
@@
Since DMFT is a local approximation, careful consideration must be given to the construction of the impurity problem. The two Cu sites experience distinct local crystal fields. Including all three Wannier functions in a single impurity problem would mix the two independent Cu ions via electronic correlations, potentially obscuring their distinct physical characters. A more appropriate approach is to define two separate impurity problems, one for each Cu site. In the following, we examine both schemes and compare their results in Fig. 2.
@@
In the first scheme, we treat all three Wannier functions within a single impurity problem. In addition to the intra-site couplings between the two degenerate orbitals on Cu-1, the $A_{1g}$ state on Cu-2 is also coupled to Cu-1 via single-particle hybridizations and density–density Coulomb interactions, resulting in an effective three-orbital impurity model. We solved this model for two interaction strengths, $U=1.0$ and $2.0$ eV, with $\beta=30$ and $J=0.1$ eV. As shown in Fig. 2(a), the resulting correlated spectral function exhibits broad features spanning from 1.5-1.5 eV to $1.0$ eV. The imaginary parts of the self-energies for the two Cu sites show distinct behaviors. The two $E_{g}$ Wannier states on Cu-1 remain nearly degenerate (red line with circles) and exhibit a smaller self-energy than that of the $A_{1g}$ state on Cu-2 (blue line with squares), indicating weaker band renormalization for Cu-1. However, increasing $U$ to $2.0$ eV lifts the degeneracy of the $E_{g}$ states, resulting in two different self-energies for Cu-1. In this unified impurity treatment, electronic correlations enhance the single-particle hybridization among the three Wannier states, leading to an unreasonable breaking of the $E_{g}$ symmetry.
@@
A more physically sound DMFT framework assigns the three Wannier states to separate impurity problems based on their spatial localization. In the second scheme, we therefore treat the two $E_{g}$ states on Cu-1 and the $A_{1g}$ state on Cu-2 in two independent impurity problems, solving for their respective self-energies separately. Both self-energies are incorporated simultaneously into the DMFT self-consistency loop and the DFT+DMFT upfolding/downfolding procedure. The resulting spectral functions and self-energies are presented in Fig. 2(c) and (d). First, similar to the unified treatment, the system remains metallic across the parameter range studied. However, the self-energy now preserves the expected symmetry: the $E_{g}$ states remain nearly degenerate and distinct from the $A_{1g}$ state at both interaction strengths. Second, we observe that the integer-filled $E_{g}$ states are more sensitive to electronic correlations. As $U$ increases from 1 to 2 eV, the self-energy of the $E_{g}$ states evolves from metallic toward more insulating behavior, while that of the $A_{1g}$ state retains its metallic character. This stark contrast arises from the difference in orbital fillings: the integer-filled $E_{g}$ states on Cu-1 experience stronger band renormalization, whereas the partially filled $A_{1g}$ state remains relatively weakly correlated. Consequently, Y3Cu2Sb3O14 exhibits a site-selective band renormalization, where bands derived from Cu-1 orbitals are more strongly correlated than those from Cu-2. The observation of site-selective band renormalization in Y3Cu2Sb3O14 may offer a promising explanation to the two-step spin condensation observed in experiment, where the charge degree of Cu-2 is less correlated and only freeze at much lower temperature, inhibiting the condensation of spins around Cu-2.
@@
This site-selective behavior is reminiscent of the extensively studied rare-earth nickelates RRNiO3 (with $R=$ Sm, Eu, Y, Lu), in which Ni occupies two distinct sublattices characterized by different Ni–O bond lengths [63]. In those materials, Ni ions in long-bond octahedra adopt a nominal $d^{8}$ configuration, whereas those in short-bond octahedra exhibit a $d^{8}\underline{L}^{2}$ configuration, where two ligand holes compensate two electrons in the Ni $e_{g}$ orbitals. This leads to inequivalent Ni sites and differential band renormalization under electronic correlations [64, 65, 66, 67]. In Y3Cu2Sb3O14, the situation is even more striking: the two Cu2+ ions experience completely inverted crystal fields, giving rise to distinct correlation responses in the two Cu sites.